OpenStudy (anonymous):

What is the area of a circle with radius 14 units, to the nearest hundredth?

pooja195 (pooja195):

\[\huge~\rm~A=~\pi~r^2 \] \[\huge~\rm~A=(3.14) ~(14)^2 \]

OpenStudy (anonymous):

196*3.14 = 615.44 ?

pooja195 (pooja195):

Are you using the google calculator?

OpenStudy (anonymous):

No I'm using my tablet calculator the Google one doesn't give me the correct answers that's why I came to here

pooja195 (pooja195):

I thought you were using it because your answer is still incorrect :P lets do this step by step \[\huge~\rm~14^2=14 \times 14 =? \]

OpenStudy (anonymous):

196

OpenStudy (k12girlohva):

Hey @reymishia Welcome to OpenStudy if i have not said that So Welcome;3 ^_^:)

pooja195 (pooja195):

Good, \[\huge~\rm~196 \times 3.14=? \]

OpenStudy (anonymous):

Thank you @k12girlohva

OpenStudy (anonymous):

I'm still getting 615.44 @pooja195

pooja195 (pooja195):

Oh right my bad that was right! I'm so sorry! >.< Welcome to openstudy! :-) Sorry for wasting your time!

OpenStudy (anonymous):

Thank you and it's okay you we're big help still @pooja165

pooja195 (pooja195):

^_^ \(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Huge\bf\color{#FF0000}W\color{#FF4900}e\color{#FF9200}l\color{#FFDB00}c\color{#FFff00}o\color{#B6ff00}m\color{#6Dff00}e\color{#24ff00}~\color{#00ff00}t\color{#00ff49}o\color{#00ff92}~\color{#00ffDB}O\color{#00ffff}p\color{#00DBff}e\color{#0092ff}n\color{#0049ff}S\color{#0000ff}t\color{#2400ff}u\color{#6D00ff}d\color{#B600ff}y\color{#FF00ff}!\)\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Large\bf\color{Lime}{Please,\ don't\ hesitate\ to\ ask\ questions!}\)\(\Large\bf\color{magenta}{We're\ here\ to\ help!}\Large\color{magenta}{\ddot\smile}\)

OpenStudy (anonymous):

Another thing I'm still trying figure out why i am not able to draw or attach anything onto here. I was told to use something else instead of my phone I tried using my tablet and I still wasn't able. @pooja165

pooja195 (pooja195):

I don't think you can draw with a tablet is it an android? And im pooja195 :P

OpenStudy (anonymous):

Yes. Another question What is the volume of a sphere that has a radius of 11.5 units? @pooja165

pooja195 (pooja195):

We currently don't have a android app it's on it's way for now you'll need to stick with writing :/ (typing:P) \[\huge~\rm~ V=\frac{ 4 }{ 3 }~\pi~r^3\] \[\huge~\rm~ V=\frac{ 4 }{ 3}(3.14) ~(11.5)^3\]

OpenStudy (anonymous):

4, 775.5475 ?

pooja195 (pooja195):

11.5^3=?

OpenStudy (anonymous):

1,520.875

pooja195 (pooja195):

Good 1520.875 x 3.14=?

OpenStudy (anonymous):

I still got 4775.5475

pooja195 (pooja195):

Good now multiply that by 4/3 \[\huge~\rm~4775.5475 \times \frac{ 4 }{ 3}\]

OpenStudy (anonymous):

6367.396

OpenStudy (anonymous):

? @pooja165

pooja195 (pooja195):

Yes perfect :)